OpenStudy (anonymous):

solve 1/2 abs(10-2x)< or = 1

OpenStudy (kc_kennylau):

\[\frac12|10-2x|\le1\]

OpenStudy (kc_kennylau):

Multiply both sides by 2 first

OpenStudy (anonymous):

So that would be \[\left| 10-2x \right| \le 2\] right?

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

Note that |y|<x means that -x<y<x

OpenStudy (kc_kennylau):

Now do the same to the question

OpenStudy (anonymous):

is it \[\left| x \le9/2 or x \ge11/2 \right|\]?

OpenStudy (kc_kennylau):

nope, please do it step by step

OpenStudy (anonymous):

\[\left| -11/2 orx \ge-9/2 \right|\]

OpenStudy (kc_kennylau):

nope don't guess and do it step by step

OpenStudy (kc_kennylau):

The answer can't help you learn

OpenStudy (kc_kennylau):

I can just tell you the answer is 4<x<6 but you won't learn if you don't know why

OpenStudy (anonymous):

how did you get that?

OpenStudy (kc_kennylau):

Again, can you do it step by step

OpenStudy (kc_kennylau):

\[\frac12|10-2x|\le1\]\[|10-2x|\le2\]\[-2\le10-2x\le2\]\[2\ge2x-10\ge-2\]\[12\ge2x\ge8\]\[6\ge x\ge4\]